hello i am having trouble trying to solve radicals with letters in them.(i know there is mathmatical word for letters i just forgot) here is one of the questions.
the square root of 360x^5y^4
thank you

Answer by scott8148(5880)

remember sqrt(ab)=sqrt(a)*sqrt(b) and set up whole roots by factoring

sqrt(360x^5y^4)=sqrt(6^2*10*(x^2)^2*x*(y^2)^2)=6x^2y^2(sqrt(10x))